OpenStudy (anonymous):

Pleaseeeeee help A certain plane flying with the wind travels 540 miles in 2 hr. later flying against the same wind the plane travels 690 mi in 3 hr. find the speed of the plan in still air and the speed if the wind?

OpenStudy (anonymous):

r*t = d speed is r+w for 2hrs and r-w for 3 hrs. you end up with two equations in two unknowns. I have to have your permission to give you the answer. Do you want only the answer now?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

Could you help me with the two equations

OpenStudy (anonymous):

\[\{2(r+w)=540,3(r-w)=690\}\]\[\{r=250\text{ mph},w=20\text { mph}\} \]where r is the speed of the plane and w is the wind speed.

OpenStudy (anonymous):

Thank you soooo much

OpenStudy (anonymous):

Your welcome.